OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (1, 9) and (–1, 11)?

OpenStudy (aroub):

slope = 11-9/ -1-1 = 2/-2 =-1 y-y1=m(x-x1) , m=-1 , (1,9) y-9=-1(x-1)

OpenStudy (anonymous):

y – 9 = 1(x + 1) y + 9 = 1(x + 1) y – 9 = –(x – 1) y + 9 = –(x – 1)

OpenStudy (aroub):

y-9=-(x-1)